OpenStudy (anonymous):

Write an equation for the line that passes through (0, -2) and whose slope is ½. y = 1/2x + 2 y = -1/2x -2 y = 1/2x – 2 y = -2x + ½

OpenStudy (cruffo):

Do you remember slope-intercept form of a line: y = mx + b.

OpenStudy (cruffo):

Here m = 1/2 and b = -2

OpenStudy (anonymous):

soooooo what answer is it?